OpenStudy (anonymous):

Evaluation of limit as n approaches infinity of (2*n^2)/(2^n)

OpenStudy (anonymous):

I need some way to show that the answer to this is 0, I'm just not sure how

ganeshie8 (ganeshie8):

apply L'hopital 2 times

ganeshie8 (ganeshie8):

nothing much happens to the denominator but the numerator reduces to a constant

ganeshie8 (ganeshie8):

\[\large \lim\limits_{n\to \infty} \dfrac{c}{2^n} = 0\]

OpenStudy (anonymous):

Not sure why I didn't see that, but thank you for pointing that out. Now would there be an easy way to do that with basically the same equation: \[ \frac{ k*n^k }{ 2^n }\] where k is an arbitrary natural number?

OpenStudy (anonymous):

or would it just be a matter of using L'Hopital's rule k number of times?

ganeshie8 (ganeshie8):

same thing, if you wait long enough, the exponential overtakes ANY polynomial

ganeshie8 (ganeshie8):

\[\large \lim\limits_{n\to \infty}\dfrac{n^{1000000000000000}}{2^n} = 0\]

ganeshie8 (ganeshie8):

and yes apply L'opital a trillion times :)

OpenStudy (anonymous):

Okay. thank you very much. Not sure why I didn't realize that, but thank you for the help!

ganeshie8 (ganeshie8):

np :) can you think of any other way to prove this without touching L'opital ?

OpenStudy (anonymous):

I know there's a way but isn't it a lot more painful then L'Hopital's, or am I thinking something else?

ganeshie8 (ganeshie8):

i remember a prof saying about this : you can avoid applying L'opital trillion times if you are clever enough.. but he didn't show the other method... i also gave up after trying for a while

ganeshie8 (ganeshie8):

from his talk, it appears the other way is more intuitive and fast let me pull up that video

OpenStudy (anonymous):

I know there are shortcuts if certain polynomials are the same, but otherwise a lot of the higher level limit work I did wasn't too fun

OpenStudy (anonymous):

I got a question about limits, away from your original question, sorry to intrude if I am, but \[\lim_{x \rightarrow {\infty}}~~~ 1^\infty = 1\] I know this is an intermediate form, but why can't we just say it = 1? What was the reason for this?

OpenStudy (anonymous):

The same old approaching from left and right?

OpenStudy (anonymous):

in the context of my question or just in general?

OpenStudy (anonymous):

General I guess

OpenStudy (anonymous):

My guess is that there isn't a particular reason, it's just to show that it is still changing and the problem isn't completely finished yet. Just a guess

OpenStudy (anonymous):

It's been troubling me, maybe have to take multiple limits?

ganeshie8 (ganeshie8):

the classic example is \[\large \lim_{n \to \infty} \left(1 + \frac{1}{n}\right)^n = e \ne 1\] even though \(\large \lim\limits_{n\to\infty}(1+\frac{1}{n}) = 1 \) and \(\large \lim\limits_{n\to\infty }n = \infty \)

ganeshie8 (ganeshie8):

@goalie2012 watch below video between `44-46` minutes : http://ocw.mit.edu/courses/mathematics/18-03-differential-equations-spring-2010/video-lectures/lecture-19-introduction-to-the-laplace-transform/ he says we can do it by applying L'opital only once but im not able to see how thats possible :/

OpenStudy (anonymous):

I don't either unless you have to move a few things around after you use the rule once? or theres a different rule you can use after that? I'm not sure.

ganeshie8 (ganeshie8):

yeah its been a mystery for me from the day i took that course haha! hope he was not thinking about taylor series >.<

OpenStudy (anonymous):

I agree. Those aren't fun at all... Might have to show that to a couple people to see what they think